# Evidence Without Injustice: A New Counterfactual Test for Fair Algorithms

### v1

@@ Abstract.
The growing philosophical literature on algorithmic fairness has examined statistical criteria such as equalized odds and calibration, causal and counterfactual approaches, and the role of structural and compounding injustices. Yet an important dimension has been overlooked: whether the evidential value of an algorithmic output itself depends on structural injustice. Our paradigmatic pair of examples contrasts a predictive policing algorithm, which relies on historical crime data, with a camera-based system that records ongoing offenses, both designed to guide police deployment. In evaluating the moral acceptability of acting on a piece of evidence, we must ask not only whether the evidence is probative in the actual world, but also whether it would remain probative in nearby worlds without the relevant injustices. The predictive policing algorithm fails this test, but the camera-based system passes it. When evidence fails the test, it is morally problematic to use it punitively, more so than evidence that passes the test.
@@ 1. Introduction
The realization that algorithmic error will fall disproportionately on minorities is likely to elicit different responses. Conservatives will point out that both algorithms track where crime is happening. They direct police to minority neighborhoods because that is where violent crime is concentrated, whether measured through historical crime data or camera recordings. The disparities in error rates across neighborhoods with different racial compositions are an inevitable, albeit unfortunate, byproduct of crime patterns. Progressives will counter that this explanation does not say why crime clusters in this way. True, areas with larger minority populations have higher rates of violent crime. [^1] By some estimates, the rate of violent crime in Black neighborhoods is, on average, five times higher than in White neighborhoods (Peterson and Krivo,, 2010). But this clustering of crime is caused by historical processes that are morally suspect, such as segregation, redlining, and disinvestment, along with their negative effects on people’s well-being today, such as poverty, poor health and crime victimization. [^2] For a causal analysis of the link between redlining and lower home ownership rates and property values, see (Aaronson et al.,, 2021). Other studies show a robust correlation between historically redlined census tracts to higher rates of negative health outcomes (Noelke et al.,, 2022). Longitudinal studies find that residents who grew up in redlined areas face a higher probability of violent victimization in adulthood, even after accounting for individual characteristics (Poulson et al.,, 2021; Testa et al.,, 2024). They also face higher probability of being killed by police (Mitchell and Chihaya,, 2022). Call these historical processes and their negative effects today structural injustices. [^3] According to a classic definition, structural injustice exists when ‘social processes put large categories of persons under systematic threat of domination or deprivation of the means to develop and exercise their capacities’ (Young,, 2011, p. 52).
Once we acknowledge that the two algorithms operate against a backdrop of structural injustice—or perhaps despite that—our moral intuitions about their fairness start to diverge. For one thing, using camera footage to locate crime or decide where to dispatch officers strikes most of us as legitimate, even if the crime itself can be traced to structural injustice. In contrast, we find it troubling when police rely on historical statistical patterns, especially patterns that are due to unjust inequalities. The two algorithms—we will argue—are not on a par from the point of view fairness. A moral difference separates them, even though both track crime patterns; both distribute algorithmic harms unevenly across demographic groups; and both operate under conditions of structural injustice. [^4] Some will also argue that both algorithms benefit minority residents the most, since increased police presence reduces crime rates and crime victimization (Risse and Zeckhauser,, 2004).
To be clear, the moral difference between the two algorithms cannot lie in whether or not they use a race-based characteristic to make assessments about the locations of crimes. It is true that a predictive policing tool that explicitly directed patrols to predominantly Black neighborhoods would be unacceptable. But the predictive policing algorithm under consideration relies on race only indirectly. It relies on crime data that is correlated with race. In contrast, the camera algorithm does use race explicitly. If footage captured “a Black man fleeing the scene,” the algorithm would incorporate that detail, along with other identifying features, to locate possible suspects. And yet this explicit use of race does not strike us as objectionable. [^5] Similarly, when a witness reports they saw a Black man near the crime scene, the identifying description ‘Black man’ is considered useful information not to be ignored (Banks,, 2001). Paradoxically, the algorithm that explicitly uses racial descriptors feels less morally troubling.
As we will see, the comparison between the two algorithms exposes a gap in current theories of algorithmic fairness. They cannot explain—or explain away—why the two algorithms seem morally different (Section 2). We think the key lies in an overlooked criterion we call the Counterfactual Independence Principle (CIP): whether the evidential basis of an algorithmic output would retain its probative value in a world without structural injustice (Sections 3). We apply this principle to both criminal justice and the health care domain (Section 4).
@@
Notice that both algorithms first forecast crime locations and, as a second step, recommend punitive responses such as increasing police presence. An algorithmic forecast about crime locations could also be used to guide supportive responses, such as modifying the built environment to block criminogenic conditions. For clarity of analysis, however, we focus first on punitive responses and only consider supportive responses at the end of the paper.
@@ 3.4. Isolating unjust social processes
Another way to push the objection is to claim that we cannot know what would happen in a world without structural injustice. How can we assess whether evidence E would still support outcome $O$ if we removed redlining, racial segregation, or concentrated poverty? Since these structural conditions have shaped the world so deeply, imagining their absence will be speculative. If so, the CIP test would rest on shaky epistemic ground. It would ask us to evaluate evidentiary relations in worlds we cannot empirically access with confidence.

### v2

@@ Abstract.
The growing philosophical literature on algorithmic fairness has examined statistical criteria such as equalized odds and calibration, causal and counterfactual approaches, and the role of structural and compounding injustices. Yet an important dimension has been overlooked: whether the evidential value of an algorithmic output itself depends on structural injustice. We contrast a predictive policing algorithm, which relies on historical crime data, with a camera-based system that records ongoing offenses, where both are designed to guide police deployment. In evaluating the moral acceptability of acting on a piece of evidence, we must ask not only whether the evidence is probative in the actual world, but also whether it would remain probative in nearby worlds without the relevant injustices. The predictive policing algorithm fails this test, but the camera-based system passes it. When evidence fails the test, it is morally problematic to use it punitively, more so than evidence that passes the test.
@@
Disclaimer on AI use: The authors used Claude 3.7 in the early stages of this project to help in the formulation of the Counterfactual Independence Principle presented in this paper. The model provided preliminary suggestions pertaining to the principle’s articulation, case studies, and the structure of the argument. Screenshots of the initial output are on file with the authors. The AI-generated material was then analyzed, critiqued, and extensively revised by the authors throughout the writing process.
@@ 1. Introduction
The realization that algorithmic error will fall disproportionately on minorities is likely to elicit different responses. Conservatives will point out that both algorithms track where crime is happening. They direct police to minority neighborhoods because that is where violent crime is concentrated, whether measured through historical crime data or camera recordings. The disparities in error rates across neighborhoods with different racial compositions are an inevitable, albeit unfortunate, byproduct of crime patterns. Progressives will counter that this explanation does not say why crime clusters in this way. True, areas with larger minority populations have higher rates of violent crime. [^1] By some estimates, the rate of violent crime in Black neighborhoods is, on average, five times higher than in White neighborhoods (Peterson and Krivo,, 2010). But this clustering of crime is caused by historical processes that are morally suspect, such as segregation, redlining, and disinvestment, along with their negative effects on people’s well-being today, such as poverty, poor health and crime victimization. [^2] For a causal analysis of the link between redlining and lower home ownership rates and property values, see (Aaronson et al.,, 2021). Other studies show a robust correlation between redlined census tracts and worse health outcomes (Noelke et al.,, 2022). Longitudinal studies find that residents who grew up in redlined areas are at greater risk of violent victimization in adulthood, even after accounting for individual characteristics (Poulson et al.,, 2021; Testa et al.,, 2024). They are also at higher risk of being killed by police (Mitchell and Chihaya,, 2022). Call these historical processes and their negative effects today structural injustices. [^3] According to a classic definition, structural injustice exists when ‘social processes put large categories of persons under systematic threat of domination or deprivation of the means to develop and exercise their capacities’ (Young,, 2011, p. 52).
Once we acknowledge that the two algorithms operate against a backdrop of structural injustice—or perhaps despite that—our moral intuitions about their fairness start to diverge. Using camera footage to locate crime or decide where to dispatch officers does not seem unfair. Even if the crime itself can be traced to structural injustice, the camera merely reports an ongoing or very recent event, and the system recommends responding to it by dispatching police patrols to the scene. To be sure, uneven placement of cameras across a city would allocate the burdens of surveillance unevenly and thereby raise fairness concerns. But the hypothetical we started out with assumed that cameras were evenly distributed. The predictive algorithm, by contrast, raises a more immediate fairness worry. It proactively sends police to areas where crime is expected to occur, and minority residents are more likely to bear the costs of unnecessary stops, heightened scrutiny, or excessive force. This disparity is particularly troubling when the algorithmic forecasts rely on data about historical crime patterns that have been shaped by structural injustice.
Admittedly, it is difficult to say exactly how, or even whether, the two algorithms differ from the standpoint of fairness. As seen earlier, both track crime patterns; both distribute algorithmic harms unevenly across demographic groups; and both operate under conditions of structural injustice. [^4] Some will argue that both algorithms ultimately benefit minority residents, since increased police presence can reduce crime rates and victimization (Risse and Zeckhauser,, 2004). In fact, the comparison between the two algorithms exposes a gap in current theories of algorithmic fairness. We will argue that prevailing theories—whether statistical, counterfactual, causal, or focused on compounding injustice and predictive justice—cannot fully explain, or explain away, why the two algorithms seem morally different (Section 2). We think that the difference lies in how structural injustice affects the evidential value of the data on which each algorithm relies. For the predictive algorithm, the data have evidential value only in a world shaped by structural injustice. The same does not hold for the camera-based algorithm. That explains why the latter algorithm feels less unfair than the former (Section 3). We then extend this analysis beyond criminal justice to the health care domain (Section 4).
@@
A couple of observations before we begin. First, we do not claim that the camera-based algorithm is morally preferable in all respects. We only claim that, from the standpoint of fairness, it is less problematic than the predictive one. Our argument does not address separate questions about privacy, legitimacy, or community consent. Camera systems may well be more troubling on those fronts. Second, notice that both algorithms first identify crime locations, and as a second step, recommend punitive responses such as increasing police presence. But an algorithmic output about a crime location could also be used to guide supportive responses, such as modifying the built environment to block criminogenic conditions. For clarity of analysis, however, we focus first on punitive responses and only consider supportive responses at the end of the paper (Section 4.3).
@@ 3.4. Isolating unjust social processes
Another way to press the objection is to claim that we cannot know what would happen in a world without structural injustice. How can we assess whether evidence E would still support outcome $O$ if we removed redlining, racial segregation, or concentrated poverty? Since these structural conditions have shaped the world so deeply, imagining their absence will be speculative. If so, the CIP test would rest on shaky epistemic ground. It would ask us to evaluate evidentiary relations in worlds we cannot empirically access with confidence.